OpenStudy (anonymous):

find the height of a square pyramid that has a volume of 25 3/5 meters and a base with 4 meter sides

OpenStudy (zpupster):

V=a2*h/3 this is the formula for the volume of a square pyramid 25 3/5 = (4)^2*h/3 can you solve for h??